Company ABC is expected to pay $1.5 dividend next year. The dividend will then grow by...

  1. Company ABC is expected to pay $1.5 dividend next year. The dividend will then grow by 10% for another three years. After that, the dividend will be constant forever. How much would you like to pay for this stock if your required rate of return is 15%?
Answer #1

Year 1 dividend = 1.5

Year 2 dividend = 1.5 (1 + 10%) = 1.65

Year 3 dividend = 1.65 (1 + 10%) = 1.815

Year 4 dividend = 1.815 (1 + 10%) = 1.9965

Value at year 3 = D4 / required rate

Value at year 3 = 1.9965 / 0.15

Value at year 3 = 13.31

Price = 1.5 / (1 + 0.15)1 + 1.65 / (1 + 0.15)2 + 1.815 / (1 + 0.15)3 + 13.31 / (1 + 0.15)3

Price = $12.50